• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mongi

    Log entry time 07:59:17 on March29,2014

    Entry number 387972

    keyword=Owl Shift Summary

    Saturday 03/29/2014
    Owl Shift Summary


    S.L :
    T.O : M. Dlamini
    3rd : F. Georges


    00:00 ; Begining of shift, NO BEAM, Expecting CW beam anytime.
    00:15 ; Beam in Hall A, 5 micro A
    checked with empty target
    raster established and verified
    01:00; Temporaly no CW beam. MCC trying to tune beam through to beam dump, so no CW beam, for the moment.
    Target moved to empty 2 to allow MCC tune beam through hall. '
    01:38; Call from MCC, reporting a tripping beam loss monitor in the hall. Their experts are working on it. We probably won'y have CW beam for about an hour.
    03:20; Call from MCC, update on beam tuning. MCC having problems with one of
    their injectors. We probably won't have CW beam till towards the end
    of the shift.

    07:15; Call from MCC. MCC has re-started tuning beam into hall

    08:00; End of shift, no Beam